OpenStudy (anonymous):

medal for easy question Solve for x. Express the answer in simplest form. 6x-1/5=3/1

OpenStudy (anonymous):

bonjour comment allez-vous

OpenStudy (anonymous):

hello

OpenStudy (texaschic101):

6x - 1/5 = 3 ---- multiply entire equation by common denominator 5 30x - 1 = 15 --- add 1 to both sides 30x = 15 + 1 30x = 16 x = 16/30 which reduces to 8/15 check.. 6x - 1/5 = 3 6(8/15) - 1/5 = 3 48/15 - 1/5 = 3 48/15 - 3/15 = 3 45/15 = 3 3 = 3 (correct) so x = 8/15

OpenStudy (texaschic101):

bon jour

OpenStudy (texaschic101):

comment allez-vous ?

OpenStudy (anonymous):

Je suis bien, merci. Et vous?

OpenStudy (anonymous):

ENGLISH